OpenStudy (lalaly):

Find \[\large{\sum_{n=1}^{\infty}\frac{1}{n^6}}\] using parsevals identity

OpenStudy (anonymous):

hey funny bunny

OpenStudy (lalaly):

lol hi

OpenStudy (anonymous):

how is your day?

OpenStudy (anonymous):

ok i don't really know this, but i am going to make a guess

OpenStudy (lalaly):

ok go ahead:D

OpenStudy (anonymous):

my guess is to use \(f(x)=x^3\)

OpenStudy (anonymous):

and before i do this i guess we (or i) should go back and look to see how parseval uses \(f(x)=x^2\) to find \[\sum \frac{1}{n^4}=\frac{\pi^4}{90}\]

OpenStudy (lalaly):

i can find sum 1/n^2 =pi^2/3 using parsevals,,, but i was wondering do i use that with cubing it for example and do the same step or im not allowed to do that

OpenStudy (anonymous):

well that was my guess, but i am still looking over the proof of \[\sum \frac{1}{n^2}=\frac{\pi^2}{6}\] and \[sum\frac{1}{n^4}=\frac{\pi^4}{90}\] to see if i understand exactly what is going on

OpenStudy (lalaly):

i can show u the proof if u want

OpenStudy (anonymous):

looks like we find \[c_n=\int_{-\pi}^{\pi}x^3e^{-inx}dx\] yes?

OpenStudy (anonymous):

will require integration by parts i guess

OpenStudy (lalaly):

ok let f(x)=x^2 after doing all the steps to find the fourier series for f(x) \[=\frac{\pi^2}{3}+\sum_{n=1}^{\infty}\frac{4(-1)^n}{n^2}\]applying parsevals identity \[\frac{1}{\pi}\int\limits_{-\pi}^{\pi}f(x)^2dx=2a_0^2+\sum_{n=1}^{\infty}(a_n^2+b_n^2)\]\[\frac{1}{\pi} \int\limits_{-\pi}^{\pi}x^4dx=2(\frac{\pi^2}{3})^2+\sum_{n=1}^{\infty}(\frac{16}{n^4)}\]now solve for the summation

OpenStudy (lalaly):

\[-\pi<x<\pi\]

OpenStudy (anonymous):

right, but now the real question boils down to how to get the coefficients to look like \(\frac{1}{n^6}\) and that is not so clear. doesn't look like \(x^3\)does that

OpenStudy (lalaly):

i tried with x^3 but didnt work

OpenStudy (anonymous):

yeah it seems like it does not. let me look around and see what i can find

OpenStudy (anonymous):

wow this is epic stuff...beyond the limits of human intellect

OpenStudy (lalaly):

ok thanku sat

OpenStudy (anonymous):

when can i start to comprehend such logic? :P

OpenStudy (anonymous):

got a lot of followers here any ideas? @Zarkon ?

OpenStudy (zarkon):

it is not coming to me yet...been a long time.

OpenStudy (anonymous):

well i am busy googling and cannot find anything similar,but it must be out there somewhere. trick is only to come up with appropriate function i think

OpenStudy (lalaly):

what is \[\zeta(6)?\]

OpenStudy (anonymous):

1

OpenStudy (anonymous):

medal please?

OpenStudy (lalaly):

http://math.stackexchange.com/questions/115981/computing-zeta6-sum-limits-k-1-infty-frac1k6-with-fourier-series this is what i found but i dont know whats the zeta function

OpenStudy (anonymous):

\[\Pi^6\div945\]

OpenStudy (anonymous):

\[\zeta(6)=\frac{\pi^4}{945}\] which i think is your answer

OpenStudy (zarkon):

problem 8 b

OpenStudy (anonymous):

isn't it pi^6 satellite?

OpenStudy (lalaly):

nice @Zarkon thats exactly what i want but is there any method or trick i can use to figure out functions for different sums for example if iw as asked for higher powers of n ...

OpenStudy (zarkon):

I might have known some tricks a long time ago...but they have all faded away.

OpenStudy (lalaly):

ok if i want to learn it ... what should i search for lol,,

OpenStudy (zarkon):

not sure :(

OpenStudy (anonymous):

yeah i saw the \(\frac{1}{n^4}\) verson a few places

OpenStudy (lalaly):

hehe thanks anyways youve done more than enough :D @Zarkon And @satellite73 Thanku for spending this much time on it you are both awesome:D

OpenStudy (zarkon):

I have a nice little book on this stuff...I might look though it tonight and if I figure out anything I'll post it.

OpenStudy (lalaly):

i'd appreciate that:)

OpenStudy (anonymous):

page 5 has a formula for \(\zeta(2n)\) http://numbers.computation.free.fr/Constants/Miscellaneous/zetageneralities.pdf